An aqueous solution of hydrogen peroxide H2O2 is 0.610 m H2O2. What is the mole fraction of hydrogen peroxide?
Answer #1

let weight of water = 1kg = 1000gm

molality   = no of moles of solute(H2O2)/weight of solvent in kg

0.610     = no of moles of solute/1

no of moles of solute(H2O2) = 0.610moles

no moles of solvent (H2O)    = 1000/18 =55.55 moles

total no of moles      = no of moles of H2O2 + no of moles of H2O

                                  = 0.610 +55.55   = 56.165 moles

mole fraction of H2O2 = no of moles of H2O2/total no of moles

                                     = 0.610/56.165 = 0.01086 >>>>> answer